A 1100 V ,50 Hz, delta connected induction motor has a star connected slip ring rotor with a phase transformation ratio of 3.8. The rotor resistance and standstill leakage reactance are 0.012 Ω/phase and 0.25 Ω/phase respectively. Neglecting the stator impedance and magnetising current determine the rotor current at 4 % slip with slip rings shorted


1
Expert's answer
2021-08-31T23:56:13-0400

"X_r=sX_s\\
X_s=0.25\Omega\\
s=0.04\\
X_r=0.04\times{0.25}=0.01\Omega\\
Z_r=\sqrt{0.012^2+0.01^2}=0.0156\Omega\\
k=\dfrac{1}{3.8}=0.26\\
Rotor\ phase\ voltage\ at\ standstill\\
E_2=1100\times{k}=1100\times{0.26}=289.5V\\
E_r=sE_2=0.04\times{289.5}=11.58V\\
I_r=\dfrac{E_r}{Z_r}=\dfrac{11.58}{0.0156}=742.3A"Rotor current at 4% slip with slip rings shorted is 742.3A